OpenStudy (anonymous):

solve for x km/3-2x=4

OpenStudy (anonymous):

km/(3-2x)=4 so multiple both side by 3-2x km=4(3-2x) km=12-8x km-12=-8x x=-(km-12)/8

OpenStudy (anonymous):

ty